# EvoSeed: Unveiling the Threat on Deep Neural Networks with Real-World Illusions

### v1

@@ 5.1. Experimental Setup
Optimizer CMA-ES: We chose to use the Vanilla Covariance Matrix Adaptation Evolution Strategy (CMA-ES) proposed by Hansen and Auger (Hansen and Auger, 2011) to optimize the initial seed vector $z$ in order to find adversarial seed vectors $z^{\prime}$ which can generate unrestricted natural adversarial samples. We initialize CMA-ES with $\mu$ with an initial seed vector and $\sigma=1$. To limit the search by CMA-ES, we also impose an $L_{\infty}$ constraint defined by the initial seed vector on the population. We further optimize for $\tau=100$ generations with a population of $\lambda=28$ individual seed vectors $z^{\prime}$. We also set up an early finish of the algorithms if we found an individual seed vector $z^{\prime}$ in the population that could misclassify the classifier model. For our experiments, we parameterize the amount of $L_{\infty}$ constraint as $epsilon$ and experiment for three different constraint values: $0.1$, $0.2$, and $0.3$.
Metrics: We perform experiments by generating $1000$ images by randomly choosing a target class for each image. We evaluate the generated images over various metrics as described below,
a) We evaluate the image by calculating the Attack Success Rate $(ASR)$ of generated images, defined as the number of images misclassified by the classifier model $F$. It defines how likely an algorithm will generate an adversarial sample.
b) We also evaluate the quality of the image by calculating two distribution-based metrics, Fréchet Inception Distance (FID) (Heusel et al., 2017), and Inception Score (IS) (Salimans et al., 2016) and by evaluating Structural Similarity Image Metric (SSIM) (Wang et al., 2004) between the images generated by non-modified initial seed vector $z$ and the optimized adversarial seed vector $z^{\prime}$.
@@
Table 1 reports the quality of images generated using non-modified initial seed vector $z$ by the variants EDM-VP and EDM-VE $(F)$ and also reports the accuracy on different classifier models $(G)$. We observe that the images generated by the variants are high image quality and classifiable by different classifier models with over $94\%$ accuracy.
5.2. Analysis of EvoSeed over Random Search (RandSeed)
Figure 3. Exemplar adversarial samples generated using EvoSeed and RandSeed algorithms. Note that EvoSeed finds high-quality adversarial samples comparable to samples from the original CIFAR-10 dataset. In contrast, RandSeed finds low-quality, highly distorted adversarial samples with a color shift towards the pure white image.
(a) (b)                               Figure 4. Accuracy on Generated Images $x$ by the classifier model $F$ over $\tau$ generations. (a) compares the performance of EvoSeed and RandSeed, while (b) compares the performance of EvoSeed with different classifier models.
In order to compare EvoSeed with Random Search (RandSeed), Table 2 compares the performance of EvoSeed with RandSeed, a random search approach to find adversarial samples. The comparison involves evaluating EvoSeed’s potential to generate adversarial samples using various diffusion and classifier models. The results presented in Table 2 demonstrate that EvoSeed discovers more adversarial samples than Random Seed and produces higher image-quality adversarial samples. The image quality of adversarial samples is comparable to that of non-adversarial samples generated by the Conditional Diffusion Model.
The disparity in image quality between EvoSeed and RandSeed is visually depicted in Figure 3. Images generated by RandSeed exhibit low quality, marked by distortion and a noticeable color shift towards white. This suggests that employing diffusion models for a simplistic search of adversarial samples using RandSeed can yield poor-quality results. Conversely, EvoSeed generates high image-quality adversarial samples comparable to the original CIFAR-10 dataset, indicating that it can find good-quality adversarial samples without explicitly being optimized for image quality.
Efficiency in finding adversarial samples is further highlighted by EvoSeed’s superiority in utilizing $L_{2}$ Robust (Wang et al., 2023) and $L_{\infty}$ Robust (Wang et al., 2023) classifiers over Standard Non-Robust (Croce et al., 2021) and Corruptions Robust (Diffenderfer et al., 2021) classifiers. This suggests that $L_{2}$ and $L_{\infty}$ Robust models were trained on slightly shifted distributions, as evidenced by marginal changes in FID scores and IS scores of the adversarial samples. Additionally, the performance of EDM-VP and EDM-VE variants is comparable, with EDM-VP discovering slightly more adversarial samples while EDM-VE produces slightly higher image-quality adversarial samples.
To comprehend the Evolution of adversarial samples across generations, Figure 5.2 (a) illustrates how EvoSeed progressively discovers more adversarial samples with increasing generations of optimization. In contrast, RandSeed quickly finds adversarial samples within the initial generations and does not improve much later. This suggests that if the random shift can potentially generate an adversarial sample, any value of the random shift can disrupt the image generation process by the conditional diffusion model. Figure 5.2 (b) reveals the impact of the chosen classifier model $(F)$ on EvoSeed’s performance. The Corruptions Robust (Diffenderfer et al., 2021) classifier demonstrates slightly greater robustness to EvoSeed than other robust and non-robust classifier variants.
@@ 5.2. Analysis of EvoSeed over Random Search (RandSeed)
5.3. Analysis of EvoSeed over $L_{\infty}$ constraint on initial seed vector
@@
Table 3. We report Attack Success Rate (ASR), Fréchet Inception Distance (FID), Inception Score (IS), and Structural Similarity Score (SSIM) for various diffusion and classifier models to generate adversarial samples using EvoSeed with different search constraints.
Models EvoSeed with $\epsilon=0.2$ EvoSeed with $\epsilon=0.3$
Diffusion Classifier Image Evaluation Image Quality Image Evaluation Image Quality
ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ ) ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ )
EDM-VP (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 88.00% 38.44 0.86 7.77 92.80% 40.29 0.81 7.90
Corruptions Robust (Diffenderfer et al., 2021) 86.30% 40.04 0.83 7.76 92.40% 41.48 0.78 7.52
$L_{2}$ Robust (Wang et al., 2023) 94.60% 44.03 0.79 7.45 98.40% 44.88 0.76 7.34
$L_{\infty}$ Robust (Wang et al., 2023) 96.00% 43.67 0.79 7.69 98.50% 45.55 0.76 7.23
EDM-VE (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 87.80% 37.05 0.86 7.59 93.70% 38.98 0.82 7.69
Corruptions Robust (Diffenderfer et al., 2021) 87.30% 39.94 0.83 7.93 92.50% 41.22 0.79 7.81
$L_{2}$ Robust (Wang et al., 2023) 95.50% 44.15 0.80 7.70 98.00% 45.30 0.76 7.54
$L_{\infty}$ Robust (Wang et al., 2023) 96.10% 44.05 0.79 7.39 98.50% 45.09 0.76 7.40
@@
Figure 5. Accuracy on Generated Images $x$ by the classifier model $F$ over $\tau$ generations. It compares the performance of EvoSeed with different search constraints $\epsilon$.
Table 4. We report attack success rate on Standard Non-Robust Classifier (Croce et al., 2021), Corruptions Robust Classifier (Diffenderfer et al., 2021), $L_{2}$ Robust Classifier (Wang et al., 2023) and $L_{\infty}$ Robust Classifier (Wang et al., 2023) for adversarial samples generated using different diffusion and classifier models.
Models Attack Success Rate (ASR) ()normal-↑($\uparrow$)( ↑ ) on Average ASR ()normal-↑($\uparrow$)( ↑ ) on
Diffusion Classifier Standard Non-Robust (Croce et al., 2021) Corruptions Robust (Diffenderfer et al., 2021) $L_{2}$ Robust (Wang et al., 2023) $L_{\infty}$ Robust (Wang et al., 2023) 4 Classifier Models
EDM-VP (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 92.80% 17.10% 14.10% 19.40% 35.85%
Corruptions Robust (Diffenderfer et al., 2021) 41.80% 92.40% 28.20% 34.70% 49.28%
$L_{2}$ Robust (Wang et al., 2023) 33.60% 39.40% 98.40% 68.70% 60.03%
$L_{\infty}$ Robust (Wang et al., 2023) 27.70% 27.10% 35.90% 98.50% 47.30%
EDM-VE (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 93.70% 15.80% 14.30% 18.60% 35.60%
Corruptions Robust (Diffenderfer et al., 2021) 43.50% 92.50% 29.50% 36.10% 50.40%
$L_{2}$ Robust (Wang et al., 2023) 34.20% 36.00% 98.00% 68.10% 59.08%
$L_{\infty}$ Robust (Wang et al., 2023) 27.60% 26.10% 37.50% 98.50% 47.43%
(a) (b)                               Figure 6. Confusion Matrix of Generated Adversarial samples using EvoSeed with $\epsilon$ = 0.3 and (a) Corruptions Robust Model (Diffenderfer et al., 2021), (b) $L_{2}$ Robust Model (Wang et al., 2023).
To comprehend the progression of adversarial samples across generations, Figure 5 illustrates how EvoSeed, under different search constraints, progressively discovers more adversarial samples as optimization generations increase.
Notably, as the search constraint is relaxed, the graph depicting EvoSeed’s performance becomes steeper, indicating a larger pool of potential adversarial seed vectors. The increased search space makes the identification of adversarial seed vectors more accessible.
6. Extended Analysis of Generated Natural Adversarial Samples
6.1. Analysis of Transferability of Generated Adversarial Samples
To assess the quality of adversarial samples, we evaluated the Attack Success Rate (ASR) across various classifier models. The transferability of adversarial samples generated under different conditions was examined, and the results are presented in Table 4, categorized by Standard Non-Robust (Croce et al., 2021), Corruptions Robust (Diffenderfer et al., 2021), $L_{2}$, and $L_{\infty}$ Robust (Wang et al., 2023) classifiers.
Analysis of Table 4 reveals that using the $L_{2}$ Robust classifier yields the highest quality adversarial samples, with approximately $60\%$ transferability across various scenarios. It is noteworthy that adversarial samples generated with the $L_{2}$ Robust classifier can also be misclassified by the $L_{\infty}$ Robust classifier, achieving an ASR of $68\%$.
We also note that adversarial samples generated by Standard Non-Robust (Croce et al., 2021) classifier have the least transferability, indirectly suggesting that the distribution of adversarial samples is closer to the original dataset as reported in Table 3. We also note that Corruptions Robust (Diffenderfer et al., 2021), $L_{2}$ Robust (Wang et al., 2023) classifiers, are the most robust models to EvoSeed algorithm with less than $43\%$ and $44\%$ ASR across different scenarios with adversarial samples generated using different diffusion and classifier models.
6.2. Class-wise Analysis of Generated Adversarial Samples
We examine the confusion matrix of adversarial samples to gain insights into the distribution of adversarial samples across conditioned labels $c$ and predicted labels by the classifier model $F$. This analysis is performed for adversarial samples generated using the Corruptions Robust (Diffenderfer et al., 2021) and $L_{2}$ Robust classifiers, as depicted in Figure 5.3.
The patterns in predicted labels are observed consistently across different classifiers through the insights provided by Figure Figure 5.3. For instance, images labeled as "Ship" predominantly tend to be predicted as "Airplane" by the classifiers. Furthermore, there is a consistent trend of confusion between specific pairs of classes, where images labeled as "Bird" are commonly misclassified as "Truck," and vice versa. Similarly, images labeled as "Cat" often experience misclassification as "Dog," and vice versa.
Significantly, these identified patterns persist irrespective of the particular classifier employed. This suggests that these inherent adversarial patterns exist within the training distribution of diffusion and classifier models, which EvoSeed effectively leverages. The black-box nature of EvoSeed complicates understanding whether these adversarial patterns are manipulated within the Diffusion Model $G$ or the Classifier Model $F$. However, a simultaneous analysis of Table 4 and Figure 5.3 indicates that EvoSeed predominantly exploits the Classifier Model rather than the Conditional Generation Model.
This is because, if the Diffusion Models were the focus of exploitation, the images generated by the Diffusion Model for the "Cat" condition would likely consist primarily of objects of the "Dog" class, leading to perfect misclassification by other classifiers since they would classify the image as "Dog" as opposed to the conditioned variable "Cat."
7. Conclusions
This study introduces EvoSeed, first-of-a-kind evolutionary strategy-based approach for generating natural adversarial samples. Our framework employs EvoSeed within a model-agnostic black-box setup, utilizing an auxiliary Conditional Diffusion Model, a Classifier Model, and CMA-ES to produce natural adversarial examples.
Experimental results demonstrate that EvoSeed excels in discovering high-quality adversarial samples closely aligned with the original data distribution, in contrast to random search methods that often yield distorted and unrecognizable adversarial samples. Additionally, we investigate the impact of varying the search space on EvoSeed’s performance and find that expanding the search space enhances the quantity of generated adversarial samples with minimal trade-off in image quality. We anticipate that this research will pave the way for new developments in generating natural adversarial samples, providing valuable insights into the limitations of classifier robustness.

### v2

@@ B.2 Hyperparameters for CMA-ES
We chose to use the Vanilla Covariance Matrix Adaptation Evolution Strategy (CMA-ES) proposed by Hansen and Auger [11] to optimize the initial seed vector $z$ to find adversarial initial seed vectors $z^{\prime}$, which can generate natural adversarial samples. We initialize CMA-ES with $\mu$ with an initial seed vector and $\sigma=1$. To limit the search by CMA-ES, we also impose an $L_{\infty}$ constraint on the population defined by the initial seed vector. We further optimize for $\tau=100$ generations with a population of $\lambda$ individual seed vectors $z^{\prime}$. We also set up an early finish of the algorithms if we found an individual seed vector $z^{\prime}$ in the population that could misclassify the classifier model. For our experiments, we defined the $\lambda$ as $(4+3*log(n))$ [11], where $n$ is a total number of parameters optimized for the initial seed vector. We also parameterize the amount of $L_{\infty}$ constraint as $\epsilon$ and use one of the following values for quantitative analysis: $0.1$, $0.2$, and $0.3$, while for qualitative analysis we use $\epsilon=0.5$.
B.3 Checking compatibility of Conditional Diffusion Model $G$ and Classifier Model $F$
Table 5: Metric values for images generated by EDM-VP, EDM-VE, and EDM-ADM variants of diffusion models for randomly sampled initial seed vector.
Metrics EDM-VP [21] EDM-VE [21]
FID [25] 4.18 4.15
Clip-IQA [26] 0.3543 0.3542
Accuracy on Standard Non-Robust [22] 95.80% 95.54%
Accuracy on Corruptions Robust [23] 96.32% 96.53%
Accuracy on $L_{2}$ Robust [24] 96.10% 95.57%
Accuracy on $L_{\infty}$ Robust [24] 93.30% 92.25%
@@ B.3 Checking compatibility of Conditional Diffusion Model $G$ and Classifier Model $F$
Table 5 reports the quality of images generated using randomly sampled initial seed vector $z$ by the variants EDM-VP and EDM-VE $(F)$ and also reports the accuracy on different classifier models $(G)$. We observe that the images generated by the variants are high image quality and classifiable by different classifier models with over $93\%$ accuracy.
B.4 Compute Resources
Table 6: Memory Requirements for Various Models Evaluated.
Model For $1$ image For $\lambda$ images
Conditional Diffusion Models $G$
SDXL-Turbo [12] 9.30 GiB 50.58 GiB
SD-Turbo [12] 3.92 GiB 32.08 GiB
PhotoReal 2.0 [13] 5.20 GiB 64.27 GiB
EDM-VP [21] 0.92 GiB 13.16 GiB
EDM-VE [21] 0.92 GiB 13.16 GiB
Classifier Models $F$
ResNet-50 [15] 0.97 GiB 3.58 GiB
ViT-L/14 [14] 3.51 GiB 48.49 GiB
Standard Non-Robust [22] 1.24 GiB 1.24 GiB
Corruptions Robust [23] 3.18 GiB 3.18 GiB
$L_{2}$ Robust [24] 5.37 GiB 5.37 GiB
$L_{\infty}$ Robust [24] 5.37 GiB 5.37 GiB
DeepFace [18] CPU CPU
Q16 [16] 1.76 GiB 9.40 GiB
NudeNet-v2 [17] CPU CPU
For the quantitative analysis, we use a single NVIDIA GeForce RTX3090 24GiB GPU, and for the qualitative analysis, we use a single NVIDIA A100 80GiB GPU. We list the GPU requirements for the different models evaluated in the experiments in Table 6.
Appendix C Comparison with Random Search
C.1 RandSeed - Random Search on Initial Seed Vector to Generate Adversarial Samples
Algorithm 2 RandSeed - Random Search on Initial Seed Vector based on Random Shift proposed by Poyuan et al. [10]
2:Initialize: $z\leftarrow\mathcal{N}(0,I)$
3:for gen in $\{1\dots\tau\}$ do
4:     for i in $\{1\dots\lambda\}$ do
5:         $\eta_{i}\sim\mathcal{U}(-\epsilon,\epsilon)$
7:         GeneratedImage $\leftarrow G(individual,c)$blacktriangleright Generate the image using $G$
8:         logits $\leftarrow F(GeneratedImage)$blacktriangleright Evaluate the image using $F$
9:         if $argmax(logits)\neq c$ then
[…]
11:         end if
12:     end for
13:end for
Based on the definition of generating adversarial sample as defined in Equation 2. We can define a random search based on the Random Shift of the initial seed vector proposed by Poyuan et al. [10]. The random shift on the initial seed vector is defined as,
$z^{\prime}=z+\mathcal{U}(-\epsilon,\epsilon)$ (9)
which incorporates sampling from a uniform distribution within the range of $-\epsilon$ to $\epsilon$ Using this random shift, we can search for an adversarial sample. We present the pseudocode for the RandSeed in the Algorithm 2.
@@ C.1 RandSeed - Random Search on Initial Seed Vector to Generate Adversarial Samples
C.2 Analysis of RandSeed over $L_{\infty}$ constraint on initial seed vector
@@
Table 7: We report Attack Success Rate (ASR), Fréchet Inception Distance (FID), Inception Score (IS), and Structural Similarity Score (SSIM) for various diffusion and classifier models to generate adversarial samples using RandSeed with $\epsilon=0.1$ as search constraint.
Diffusion Model $G$ Classifier Model $F$ Image Evaluation Image Quality
ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ )
EDM-VP [21] Standard Non-Robust [22] 57.10% 126.94 0.25 3.72
Corruptions Robust [23] 51.50% 124.36 0.25 3.81
$L_{2}$ Robust [24] 47.60% 125.44 0.24 3.85
$L_{\infty}$ Robust [24] 49.60% 124.03 0.25 3.75
EDM-VE [21] Standard Non-Robust [22] 50.20% 112.39 0.28 4.51
Corruptions Robust [23] 42.90% 111.93 0.28 4.42
$L_{2}$ Robust [24] 42.70% 112.51 0.28 4.40
$L_{\infty}$ Robust [24] 40.30% 109.92 0.28 4.45
In order to compare EvoSeed with Random Search (RandSeed), Table 7 presents the performance of RandSeed, a random search approach to find adversarial samples. We generate $1000$ images with Random Seed for evaluation. The comparison involves evaluating EvoSeed’s potential to generate adversarial samples using various diffusion and classifier models. The results presented in Table 7 demonstrate that EvoSeed discovers more adversarial samples than Random Seed and produces higher image-quality adversarial samples. The image quality of adversarial samples is comparable to that of non-adversarial samples generated by the Conditional Diffusion Model.
C.3 Analysis of Images generated by EvoSeed compared to Random Search (RandSeed)
Figure 8: Exemplar adversarial samples generated using EvoSeed and RandSeed algorithms. Note that EvoSeed finds high-quality adversarial samples comparable to samples from the original CIFAR-10 dataset. In contrast, RandSeed finds low-quality, highly distorted adversarial samples with a color shift towards the pure white image.
The disparity in image quality between EvoSeed and RandSeed is visually depicted in Figure 8. Images generated by RandSeed exhibit low quality, marked by distortion and a noticeable color shift towards white. This suggests that employing diffusion models for a simplistic search of adversarial samples using RandSeed can yield poor-quality results. Conversely, EvoSeed generates high-image-quality adversarial samples comparable to the original CIFAR-10 dataset, indicating that it can find good-quality adversarial samples without explicitly optimizing them for image quality.
Appendix D Extended Qualitative Analysis of Adversarial Images generated using EvoSeed
D.1 Analysis of Image for Object Classification
Figure 9: We provide some exemplar adversarial images created by NanoSD [28].
We present some exemplar adversarial images in Figure 10 created by NanoSD [28] that are misclassified as reported in Table 4.
D.2 Analysis of Image for Ethnicity Classification
Figure 10: Adversarial images created with EvoSeed serve as prime examples of how to deceive a range of classifiers tailored for various tasks.
Figure 11: Adversarial images created with EvoSeed serve as prime examples of how to deceive a range of classifiers tailored for various tasks.
We present some more exemplar images where ethnicity of an individual can be misclassified in Figure 10. We also provide some more exemplar cases where gender of an individual was misaligned in the generate image with the given conditioning $c$ as shown in Figure 11.
Appendix E Extended Quantitative Analysis of Adversarial Images generated using EvoSeed
E.1 Analysis of Images Generated over the generations
EvoSeed with $\epsilon=0.1$
@@ E.1 Analysis of Images Generated over the generations
Figure 12: Accuracy on Generated Images $x$ by the classifier model $F$ over $\tau$ generations. (a) compares the performance of EvoSeed and RandSeed, while (b) compares the performance of EvoSeed with different classifier models.
(a) EvoSeed with $\epsilon=0.2$ [Uncaptioned image] (b) EvoSeed with $\epsilon=0.3$ [Uncaptioned image] (c)
Here, we analyse the EvoSeed’s performance with respect to the number of generations, as shown in Figure E.1. We observe that, for EvoSeed with $\epsilon=0.1$, the curves do not saturate suggesting that a higher number of generations to craft natural adversarial samples will further improve the attack performance.